# spacing=nonfrench

### v1

@@ 3.2 Trust Extension to Accelerators and Storage
LLM-based agents require GPU acceleration (for inference) and persistent storage, yet standard CVMs protect only CPU memory. GPU memory remains accessible to hypervisors, exposing model weights and inference results, while CVMs lack confidential storage support and suffer from I/O performance degradation. Figure 2b shows that VMs achieve 380K/247K IOPS for 4KB random reads/writes, CVMs reach 274K/207K, while bare metal storage stack achieves 629K/285K, revealing kernel mediation overhead as the primary bottleneck.
@@ 3.4 Multi-agent Orchestration and Communication
Large-scale simulations using Azure Functions traces (12 hours, 10,000 agents, 1M prompts) shows per-agent CVMs reach p50/p99 scheduling delays of 1,827.55/38,364.79 ms—2.05/22.52 higher than container-based approaches (Figure 3a), and incur p50/p99 slowdowns of 1.04/3.76—1.02/3.05 higher than container based approaches (Figure 3b). Beyond scheduling, CVMs impose substantial networking penalties through buffer copies and mandatory encryption (e.g., TLS) [cheng_tdxsurvey_2023, misono2024confidential]. Figure 3c shows CVMs demonstrate up to 44% higher inter-instance latency compared to containers.
@@ 4 Omega Overview
We present Omega, a system that deploys multiple trusted AI agents within a single CVM, is equipped with efficient resource management and communication mechanisms, enforces specified policies, and maintains verifiable audit trails.
@@ 4.1 System Overview
Figure 4 depicts the overview of Omega. It executes within a CVM that provides hardware-enforced isolation protecting agent code, LLM models, and user data from the untrusted cloud provider. Omega leverages AMD SEV-SNP’s VMPLs [snp-white-paper] to organize components across three privilege levels.
@@ 4.3 Programming Model and APIs
Omega exposes system APIs to support the agent lifecycle (Table 1). The Agent Provider API enables registration of agent images, LoRA weights, and baseline policies. The User API supports task submission, platform attestation. and result retrieval with tamper-evident proofs for retrospective auditing.
@@ 5 Trusted Agent Platform (TAP)
Omega’s TAP (Figure 5) defines the agent execution units (trustlets) and resource management primitives (Table 3) that enable the agent orchestrator (§ 6) to deploy agents in sandboxed environments. It aims to ensure (i) confidentiality and integrity of execution, complemented with capabilities for verification of the entire deployment (§ 5.1), (ii) strong isolation and sandboxing between co-located agents (§ 5.2), and (iii) secure device accesses across heterogeneous computing resources including GPUs (§ 5.3) and external storage (§ 5.4).
@@ 5.1 Trusted Monitor
CVM partitioning. Therefore, Omega leverages AMD SEV-SNP’s VMPLs [snp-white-paper] and employs an SVSM-based trusted monitor [coconut] operating at VMPL-0, while other components run at higher VMPLs (Figure 5). This privilege separation is enforced through the Reverse Map table, which tracks page ownership and access permissions per VMPL [snp-white-paper]. The trusted monitor uses the RMPADJUST instruction to manage lower VMPL permissions, ensuring that the agents or the TAP cannot access its memory or tamper with vTPM measurements.
@@ 5.2 Trustlets: Sandboxed Isolated Execution Units
Therefore, TAP provides a trustlet abstraction, which is a sandboxed execution unit specifically designed for agent operations within a CVM. Omega executes each trustlet at VMPL-2, thereby preventing it from accessing the memory of the TAP and the trusted monitor. Each trustlet maintains its own virtual address space while the TAP configures their page tables to ensure proper isolation between them. Trustlets do not have direct I/O access, including network access, and must invoke the TAP to communicate with other entities.
@@ 5.3 Confidential Accelerated Subsystem
delete_truslet(tid) Delete a trustlet.
@@ 5.4 AgentStore with Direct I/O Engine
AgentStore. The AgentStore provides a shared-log abstraction [jia2021boki, balakrishnan2012corfu] to trustlets, allowing each trustlet to concurrently access and append totally ordered log entries. This serves as a building block of agents’ long-term memory [packer2024memgpt, zhong2024memorybank], which contains consolidated knowledge accumulated throughout the agent’s lifetime (§ 6.3). AgentStore maintains one log per application, ensuring not only the confidentiality and integrity of data but also its freshness to prevent rollback attacks.
Trustlets append log entries using the append_log() API. Internally, TAP enhances each log entry with a monotonically increasing counter value provided by a trusted counter [speicher, rote] to detect rollback attacks, and encrypts the entry with an authenticated encryption cipher [wu2013aegis] to ensure its integrity and confidentiality. When reading entries from the AgentStore, TAP first decrypts the entry, validates its integrity and freshness using counter values, and returns the actual log content.
@@ 7.1 Policy Language and Compiler
Declarative policy specification. Omega introduces a declarative policy language for AI agents inspired by prior work [vahldiek2015guardat, krahn2018pesos, unnibhavi2022ironsafe]. Omega considers three distinct communication channels: agent-to-LLM (Omega API), agent-to-agent (A2A), and agent-to-tool (MCP). While agent-to-LLM calls may generate unwanted planning, actual actions occur via A2A/MCP calls. Therefore, Omega focuses on providing fine-grained control over MCP and A2A protocol actions, thereby enforcing strict control over the agent’s capabilities.
@@ 7.4 Policy Specification: Case-study
Case-study: Backdoor prevention. We consider a computer-use agent that interacts with an MCP server running locally on the user’s computer that exposes file operations (e.g., open(path) and write(file, content)).
@@ 9 Evaluation
Prototype. We implement a Omega prototype on AMD SEV-SNP CVMs and NVIDIA H100 CGPUs. The trusted monitor is based on COCONUT-SVSM [coconut]. We implement TAP using Linux and write the agent orchestrator in Python. We use llama-cpp [llama_cpp] as our LLM service and extend it to support shared memory communication. The current prototype lacks VMPL-based isolation; therefore, we perform our evaluation without VMPL isolation and trusted boot. Complementarily, we estimate the overhead using a microbenchmark (§ 9.2).
@@ 1 CVM containing agents w/ $C$ GPU for inference
Workload. For our end-to-end and policy evaluation, we use the WebArena [zhou2023webarena] and MCPSecBench [yang2025mcpsecbench] benchmarks. For our inference analysis, we run synthetic workloads using the ShareGPT [shareGPT] dataset, containing prompts and responses from chatbot conversations. Lastly, we conduct simulation-based scalability studies using Azure traces [serverless_in_the_wild, harvested_serverless].
@@ 9.2 Trusted Agent Platform (TAP)
Boot time analysis. We compare boot time without trusted boot for fair comparison. As shown in Figure 9c, the VM baseline exhibits high boot times (6.5s), with CVM being even slower due to additional management tasks (e.g., measurement of the initial state). Omega achieves lower start-up latency by eliminating the costly initialization of the VMM (QEMU), Firmware (OVMF) and Guest OS (Linux). This leads to a 65.2 and 181.2 improvement over the VM and CVM cases. Omega experiences a 3.3 slower bootup time than the Container baseline due to the extra initialization of the agent runtime (e.g., shared-memory channel setup).
Attestation time. We set up an attestation service and communicate with it via HTTPS. For data measurements (e.g., model, inference engine), we use SHA-384. We attest the GPU using NVIDIA’s SDK [nvidia-attestation-sdk] and the CVM using snpguest [snpguest]. Attestation latencies are averaged over 100 runs.
@@ 9.3 Agent Orchestration
Variant
Register
request (ms)
Inference
Return
result (ms)
Overall
Container 2.63 15.38 2.37 20.38
VM 5.05 16.14 1.49 22.67
CVM 3.94 26.09 1.75 31.78
Omega 2.24 28.67 0.43 31.33
Table 8: TTFT breakdown (Llama 3.1 8B).
@@ 10 Related Work
OS-level compartmentalization. Prior work explores OS compartmentalization for fine-grained isolation. Some leverage hardware-assisted page tagging (e.g., MPK [intel_mpk]) for intra-process or intra-kernel isolation [lefeuvre_flexos_2022, cubicleOS, li_iso-unik_2020, zhang_erebor_2025, sung2020unimpk, hedayati_hodor_2019, vahldiek2019erim]. For SEV-SNP, multiple studies utilize VMPL [sev-snp-abi, sev-snp-vmpl] to create isolated environments within CVMs [ge_hecate_2022, mai2023honeycomb, veil_ahmad_2024, schwarz2025seven, mei2024cabin, chen2024cpc, wang_nestedsgx_2025]. Omega leverages VMPL to create strict sandboxed environments tailored for multi-tenant trusted agent execution.
@@ Appendix A MCPSecBench Attack Details
Policy: Mitigations require restricting access either to the sensitive files or to the exfiltration channels.
@@
Policy: Our policy language can specify the number of times a tool is allowed to be called.
@@
Policy: The files allowed to be accessed can be explicitly defined.
@@
Privilege escalation. In addition to the agent context, tool descriptions may also impose additional constraints on the agents. In particular, tools may impose restrictions on the types of agents allowed to call them. For example, a "show_credentials" tool may be needed for a privileged shopping agent that needs to log into the user’s account, but other agents should not have access to this function.
@@
Policy: Tools require explicit permission in policy. Therefore, if the policy does not include a rule allowing the show_credentials tool for the agent, it is considered disallowed by default.

### v2

@@ 3.2 Trust Extension to Accelerators and Storage
LLM-based agents require GPU acceleration (for inference) and persistent storage, yet standard CVMs protect only CPU memory. GPU memory remains accessible to hypervisors, exposing model weights and inference results, while CVMs lack confidential storage support and suffer from I/O performance degradation. Figure 2b shows that VMs achieve 380K/247K IOPS for 4KB random reads/writes, CVMs reach 273K/207K, while bare metal storage stack achieves 629K/285K, revealing kernel mediation overhead as the primary bottleneck.
@@ 3.4 Multi-agent Orchestration and Communication
Large-scale simulations using Azure traces (12 hours, 10,000 agents, 1M prompts) shows per-agent CVMs reach p50/p99 scheduling delays of 1,827/38,364 ms — 2/22.5 higher than container-based approaches (Figure 3a), and incur p50/p99 slowdowns of 1.04/3.76 — 1.02/3.05 higher than container based approaches (Figure 3b). Beyond scheduling, CVMs impose substantial networking penalties due to buffer copies and mandatory encryption (e.g., TLS) [cheng_tdxsurvey_2023, misono2024confidential], as demonstrated in Figure 3c.
@@ 4 Omega Overview
We present Omega, the first unified architecture for deploying trusted AI agents inside CVMs. It consolidates multiple mutually untrusted agents within a single hardware-isolated environment, enforces the execution constraints and auditability required for trusted behavior, and supports fast communication and resource sharing without weakening security.
@@ 4.1 System Overview
Figure 4 depicts the overview of Omega. It executes within a CVM that provides hardware-enforced isolation protecting agent code, LLM models, and user data from the untrusted cloud provider. Omega leverages AMD SEV-SNP’s VMPLs [sev-snp] to organize components across three privilege levels.
@@
This decomposition reflects a deliberate balance: we isolate low-level enforcement mechanisms in TAP, centralize control in the orchestrator, and keep policy and trust establishment orthogonal so that Omega can evolve with new workloads and hardware capabilities without redesigning the entire stack.
@@ 4.3 Programming Model and APIs
Omega’s programming model is tailored to agent-development workflows. We provide easy-to-use API calls for all the building blocks of agent application development (inference, tool interactions, communication), and extend them with attestation and policy specification capabilities.
In particular, Omega exposes system APIs to support the agent lifecycle (Table 1). The Agent Provider API enables registration of agent images, LoRA weights, and baseline policies. The User API supports task submission, platform attestation. and result retrieval with tamper-evident proofs for retrospective auditing.
@@ 5 Trusted Agent Platform (TAP)
Omega’s TAP (Figure 5) provides the low-level execution substrate that isolates untrusted agents, exposes narrow interfaces to the upper layers, and offers verifiable execution to the orchestrator and policy engine. Its design keeps efficiency-critical mechanisms in this substrate while delegating supervision and cross-principal trust enforcement to higher layers. TAP defines the agent execution units (trustlets) and resource-management primitives (Table 3) that allow the orchestrator (§ 6) to deploy agents in sandboxed environments. It ensures (i) confidentiality and integrity of execution with support for full-stack verification (§ 5.1), (ii) strong isolation between co-located agents (§ 5.2), and (iii) secure access to heterogeneous resources such as GPUs (§ 5.3) and persistent storage (§ 5.4).
@@ 5.1 Trusted Monitor
Omega introduces a Trusted Monitor to establish a narrow, verifiable security core inside the CVM. This design allows us to anchor Omega’s trust guarantees in a minimal, privileged component while keeping higher-layer mechanisms flexible and updatable.
@@
CVM partitioning. Therefore, Omega leverages AMD SEV-SNP’s VMPLs [sev-snp] and employs an SVSM-based trusted monitor [coconut] operating at VMPL-0, while other components run at higher VMPLs (Figure 5). This privilege separation is enforced through the Reverse Map table, which tracks page ownership and access permissions per VMPL [sev-snp]. The trusted monitor uses the RMPADJUST instruction to manage lower VMPL permissions, ensuring that the agents or the TAP cannot access its memory or tamper with vTPM measurements.
@@ 5.2 Trustlets: Sandboxed Isolated Execution Units
Therefore, to isolate agents from the environment, TAP provides a trustlet abstraction, which is a sandboxed execution unit specifically designed for agent operations within a CVM. Omega executes each trustlet at VMPL-2, thereby preventing it from accessing the memory of the TAP and the trusted monitor. Each trustlet maintains its own virtual address space while the TAP configures their page tables to ensure proper isolation between them. Trustlets do not have direct I/O access, including network access, and must invoke the TAP to communicate with other entities.
@@ 5.3 Confidential Accelerated Subsystem
delete_trustlet(tid) Delete a trustlet.
@@ 5.4 AgentStore with Direct I/O Engine
AgentStore. Agents consider past decisions and actions when taking new decisions. Therefore, altering the memory storing their past decisions can easily influence their present and future actions. Omega provides the AgentStore, a shared-log abstraction [jia2021boki, balakrishnan2012corfu], allowing each trustlet to concurrently access and append totally ordered log entries. This serves as a building block of agents’ long-term memory [packer2024memgpt, zhong2024memorybank], which contains consolidated knowledge accumulated throughout the agent’s lifetime (§ 6.3). AgentStore maintains one log per application, ensuring not only the confidentiality and integrity of data but also its freshness to prevent rollback attacks.
Trustlets append log entries using the append_log() API. Internally, TAP enhances each log entry with a monotonically increasing counter value provided by a trusted counter [speicher, anchor, rote] to detect rollback attacks, and encrypts the entry with an authenticated encryption cipher [wu2013aegis] to ensure its integrity and confidentiality. When reading entries from the AgentStore, TAP first decrypts the entry, validates its integrity and freshness using counter values, and returns the actual log content.
@@ 7.1 Policy Language and Compiler
Declarative policy specification. Omega introduces a declarative policy language for AI agents inspired by prior work [vahldiek2015guardat, krahn2018pesos, unnibhavi2022ironsafe]. Omega considers three distinct communication channels: agent-to-LLM (Omega API), agent-to-agent (A2A), and agent-to-tool (MCP). While agent-to-LLM calls may generate unwanted planning, actual actions occur via A2A/MCP calls. Therefore, Omega focuses on providing fine-grained control over MCP and A2A protocol actions, thereby enforcing strict control over the agent’s capabilities, and ensuring policy compliance even in the presence of agent hallucinations.
@@ 7.4 Policy Specification: Case-study
Case-study: Backdoor prevention. We assume a computer-use agent interacting with a local MCP server on the user’s device that exposes file operations (e.g., open(path) and write(file, content)).
@@ 9 Evaluation
Prototype. We implement a Omega prototype on AMD SEV-SNP CVMs and NVIDIA H100 CGPUs. The trusted monitor is based on COCONUT-SVSM [coconut]. We implement TAP using Linux and write the agent orchestrator in Python. We use llama.cpp [llama_cpp] as our LLM service and extend it to support shared memory communication. The current prototype lacks VMPL-based isolation; therefore, we perform our evaluation without VMPL isolation and trusted boot. Complementarily, we estimate the overhead using a microbenchmark (§ 9.2).
@@ 1 CVM containing agents w/ $C$ GPU for inference
Workload.
For our end-to-end and policy evaluation, we use the WebArena [zhou2023webarena] and MCPSecBench [yang2025mcpsecbench] benchmarks. For our inference analysis, we run synthetic workloads using the ShareGPT [shareGPT] dataset, containing prompts and responses from chatbot conversations. Lastly, we conduct simulation-based scalability studies using Azure traces [serverless_in_the_wild, harvested_serverless].
@@ 9.2 Trusted Agent Platform (TAP)
Boot time analysis. We compare boot time without trusted boot for fairness. As shown in Figure 9c, the VM baseline exhibits high boot times (6.5s), with CVM being even slower due to additional management tasks (e.g., measurement of the initial state). Omega achieves lower start-up latency by eliminating the costly initialization of the VMM (QEMU), Firmware (OVMF) and Guest OS (Linux). This leads to a 65.2 and 181.2 improvement over the VM and CVM cases. Omega experiences a 3.3 slower bootup time than the Container baseline due to the extra initialization of the agent runtime (e.g., shared-memory channel setup).
Attestation time. We set up an attestation service accessed via HTTPS and use SHA-384 for data measurements (e.g., model, inference engine). The GPU is attested with NVIDIA’s SDK [nvidia-attestation-sdk] and the CVM with snpguest [snpguest]. Figure 2c shows that attesting the platform, model, and agent induces a latency of 8.3 s, 15.6 s, and 1.6 s, respectively, while attesting user input/output adds 200 ms; many of these costs can be reduced by caching attestation reports. Attestation latencies are averaged over 100 runs.
@@ 9.3 Agent Orchestration
Variant
Register
request (ms)
Inference
Return
result (ms)
Overall
Container 2.63 15.38 2.37 20.38
VM 5.05 16.14 1.49 22.67
CVM 3.94 26.09 1.75 31.78
Omega 2.24 28.67 0.43 31.33
Table 9: TTFT breakdown (Llama 3.1 8B).
@@ 10 Related Work
OS-level compartmentalization. Prior work explores OS compartmentalization for fine-grained isolation. Some leverage hardware-assisted page tagging (e.g., MPK [intel_mpk]) for intra-process or intra-kernel isolation [lefeuvre_flexos_2022, cubicleOS, li_iso-unik_2020, zhang_erebor_2025, sung2020unimpk, hedayati_hodor_2019, vahldiek2019erim]. For SEV-SNP, multiple studies utilize VMPL [sev-snp-abi, sev-snp-vmpl] to create isolated environments within CVMs [ge_hecate_2022, mai2023honeycomb, veil_ahmad_2024, schwarz2025seven, mei2024cabin, chen2024cpc, wang_nestedsgx_2025, sabanic2025confidentialserverlesscomputing]. Omega leverages VMPL to create strict sandboxed environments tailored for multi-tenant trusted agent execution.
@@ Appendix A MCPSecBench Attack Details
Policy: Mitigations require restricting access either to the sensitive files or to the exfiltration channels. We can either disallow files that contain secrets (e.g., “API_keys.txt”) or only allow trusted communication partners (e.g., “123@yahoo.com”), ensuring that any sensitive data accessed by an agent cannot be sent to an external attacker.
@@
Policy: Our policy language can specify the number of times a tool is allowed to be called. Thus, if we expect a prompt to, for example, only result in the acquisition of one single item, we can explicitly specify that in the policy.
@@
Policy: The files allowed to be accessed can be explicitly defined, preventing access to arbitrary files.
@@
Privilege escalation. In addition to the agent context, tool descriptions may also impose additional constraints on the agents. In particular, tools may impose restrictions on the types of agents allowed to call them. For example, a “show_credentials” tool may be needed for a privileged shopping agent that needs to log into the user’s account, but other agents should not have access to this function.
@@
Policy: In Omega, tools require explicit permission in the policy. Therefore, if the policy does not include a rule allowing the show_credentials tool for the agent, it is considered disallowed by default.